Shift summary for Owl, 10/2/99

Ed Blucher (blucher@hep.uchicago.edu)
Sat, 2 Oct 1999 06:57:31 -0500


Shift Summary for Owl Shift, 10/2/99

On Shift: Ed Blucher, Mike Arenton

Hours of beam available at all intensities: 7.0
Hours of beam available at nominal intensity: 7.0
Hours of E799 data (special runs not included): 6.2
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: DDD Veto, tape change, resetting fscc's, kfrends

========================================================================
Runs ended this shift:
Run 14791 - BEAM_799
Run 14793 - BEAM_799 in progress

========================================================================
Accesses during shift: none

========================================================================
Overview of shift:
E799 datataking (SEM=8.5E12). A few spills with dead time because
KTeV4 didn't finish in time. This sometimes occurred in spills with
relatively low intensity. From the front-panel lights, it looked like
data transfers from the ktev4 ddds sometimes stopped for a few
seconds in the middle of a spill.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14791 0:15 Start run (needed to reset fscc's)

0:30 Notice that SOD means are low by ~100 microns.
(Hogan says that Ashkan knows about this.)
1:15 DDD Veto 3 ends run. No end of tape mark
so we load fresh tapes.
1:20 New CSQ6 constants installed (valid for 14792->)

14793 1:40 Start run

1:45 Mount UPX004 for online split.

2:20 Mount UPK006 for online split.

3:15 Mount UPH006 for online split.

4:50 Mount UPE003 for online split.

197 6:05 Huge laser rms for ~800 crystals. The pattern
looked like a pipeline problem, and was cured
after a couple of pipeline resets.

6:35 Mount UPM011 for online split.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

BobT would like to be paged if there is no beam for >= 1 hour
(day/evening).